# One-dimensional lattice random walks in a Gaussian random potential

### v2

@@ Abstract
We study random walks evolving in continuous time on a one-dimensional lattice where each site $x$ hosts a quenched random potential $U_{x}$. The potentials on different sites are independent, identically distributed Gaussian random variables. We analyze three distinct models that specify how the transition rates depend on $U_{x}$: the random-force-like model, random walks with randomized stepping times, and the Gaussian trap model. Our analysis focuses on five key disorder-dependent quantities defined for a finite chain with $N$ sites: the probability current, its reciprocal (the resistance), the splitting probability, the mean first-passage time $T_{N}$, and the diffusion coefficient $D_{N}$ in a periodic chain. By determining the moments of these random variables, we demonstrate that the probability current, resistance, and splitting probability are not self-averaging, which leads to pronounced differences between their average and typical behaviors. In contrast, $T_{N}$ and $D_{N}$ become self-averaging when $N\to\infty$, though they exhibit strong sample-to-sample fluctuations for finite $N$.
Keywords: Random walks in disordered media, self-averaging vs non-self-averaging
December 31, 2025
[…]
@@ 1 Introduction
In this paper, we revisit the long-standing problem of random walks in the presence of a quenched random Gaussian potential (Gaussian white noise), a topic that has received a significant attention in the past (see, for example, [51]). Specifically, we study dynamics of a particle that undergoes, in continuous time, a random walk between nearest-neighboring sites of a one-dimensional lattice, where each site $x$ hosts a quenched random potential $U_{x}$, which influences the local transition rates. The variables $U_{x}$ are independent and identically distributed (iid) random variables, with a common Gaussian distribution. This model can be also viewed from a somewhat different perspective. Namely, it can be considered as an effective description, in terms of the Fick-Jacob’s approximation (see, e.g., [60, 61]), of diffusion in a narrow two- or three-dimensional channel with Gaussian-distributed random wall corrugation.
@@
While the statistics of the last two quantities have been explored in previous studies and some well-established results have been obtained (see, for example, [49, 48, 51]), a comprehensive analysis of the first three variables is lacking at present. It is also important to note that all the random variables under study are formally defined in the limit $t\to\infty$. Consequently, our analysis does not address transient effects, such as aging and the ensuing sub-diffusive behavior, which have been extensively discussed in the literature on the random force and random traps models (see, for instance, [49]).
@@
The paper is outlined as follows: In Sec. 2 we present our model and describe three different dynamical scenarios. In Sec. 3 we concentrate on the statistical properties of the realization-dependent currents through a finite chain and of its reciptocal value - the resistance of a finite chain with respect to a random transport. In Sec. 4 we discuss the behavior of the realization-dependent splitting probabilities. Section 5 is devoted to the analysis of the mean first-passage times through finite intervals. Lastly, in Sec. 6 we evaluate the moments of the diffusion coefficient in period chains in the limit $N\to\infty$ for all three dynamical scenarios. In Sec. 7 we conclude with a brief recapitulation of our results. Some derivations are relegated to Appendices: in A we present the derivation of the probability currents through a finite chain with fixed disorder for all three dynamical scenarios, while in B, capitalizing on the exact result due to Derrida [4], we present exact compact expressions for the realization-dependent diffusion coefficient.
@@ 2.2 Scenario II: Random walk with randomized stepping-times
Supposing that each site of the lattice hosts a random potential $U_{x}$, we define the transition probabilities $p_{x,x\pm 1}$ in the way which resembles the definition of the transition rates in the above Model I, except that these probabilities are normalized to unity. That being, we stipulate that
@@ 3.1 The moments of resistances
The above equations imply that the resistances in Models I and II are not self-averaging in the limit $N\to\infty$. The resistance in the Model III is not self-averaging for any value of $N$. Moreover, in all three Models the relative variances $R_{\tau}$ show an unbounded growth with $\beta\sigma$, which signifies that fluctuations become more important the larger the disorder is. The behavior of $R_{\tau}$ for all three Models is depicted in Fig. 2. Below we will show that such a lack of self-averaging is essentially the boundary effect.
Figure 2: The relative variances $R_{\tau}$ of the resistances $\tau_{N}$ in the Models I (left panel), II (middle panel) and III (right panel) as functions of $\beta\sigma$. Solid curves are our analytical predictions in eqs. (28) to (30) for $N=\infty$, while the symbols depict $R_{\tau}$ evaluated by a numerical averaging $\tau_{N}$ and $\tau^{2}_{N}$ for finite values of $N$ (see the insets).
@@
It is easy to find that the first two moments of the latter probability density function are given precisely by eq. (LABEL:momtau2). Note, as well, that the probability density functions in eqs. (37) and (38) have the form of the log-normal distribution as the parental form in eq. (5), and differ from it (and between themselves) only in the pre-exponential algebraic dependence on the reduced resistance and the normalizations.
@@
– The lack of self-averaging is solely due to fluctuations of the on-site potentials at sites $x=0$ (for Models I and III) and at the sites $x=0,1$ (for Model II). The contributions of the sites in the bulk of the chain are self-averaging for Models I and II, and is constant, i.e., non-fluctuating for the Model III.
@@ 3.3 Probability currents
i.e., the typical current in these two Models is characterized by the same decreasing, super-Arrhenius function of disorder and therefore is much smaller than the averaged currents in eqs. (43). Such a strong "reduction" of the current is apparently due to the fact that in typical realizations of disorder a particle will always get localized between two sites at which sufficiently large forces point in the opposite directions. Further on, we find that in the Model III the typical current obeys
@@ 3.4 Sample-to-sample fluctuations of the resistances and currents
This probability density function is depicted in Fig. 5 for three values of the parameter $\beta\sigma$. We observe that for a relatively small disorder, i.e., for $\beta\sigma=0.1$, $P(\omega)$ is unimodal and centered at $\omega=1/2$. This signifies that for a small disorder it is indeed most likely that the values of resistances (or currents) observed in two randomly chosen samples are very close to each other. Upon an increase of $\beta\sigma$, for $\beta\sigma=1$, we encounter a different behavior when the central part of $P(\omega)$ becomes flat and $\omega=1/2$ is no longer the most probable value. In this situation, (almost) any relation between the two resistance becomes equally probable. Increasing $\beta\sigma$ further, we have that already for $\beta\sigma=1.2$ the probability density function becomes bimodal with two maxima close to $\omega=0$ and $\omega=1$. Here, $\omega=1/2$ corresponds not to the most probable but to the least probable value. This signifies that the resistance observed in two sample are most likely disproportionally different and sample-to-sample fluctuations are very significant.
@@ 4 The realization-dependent splitting probability
Focusing on the Models I and II (we do not expect a very different behavior to appear in the Model III), we consider the probability $E_{-}$ that a particle starting at position $x_{0}$ on a finite interval of length $N$ first approaches the left boundary $x=0$ without ever touching the right boundary $x=N$. This probability is explicitly given by (see, e.g., [45])
$E_{-}=\frac{\tau_{N-x_{0}}}{\tau_{N-x_{0}}+\tau_{x_{0}}}\,$, (53)
where $\tau_{N-x_{0}}$ and $\tau_{x_{0}}$ are the resistances of the intervals from the right and from the left from the starting point. These are resistances are evidently independent random variables.
We introduce a relative distance $\alpha=x_{0}/N$ to the left boundary and assume that $N$ is very large and $\alpha$ is fixed and bounded away from 0 and $1$. Using the reduced variable $\overline{\tau}=\tau_{N}/N$, we then formally rewrite eq. (53) as
$E_{-}=\frac{(N-x_{0})\overline{\tau}_{1}}{(N-x_{0})\overline{\tau}_{1}+x_{0}\overline{\tau}_{2}}=\frac{(1-\alpha)\overline{\tau}_{1}}{(1-\alpha)\overline{\tau}_{1}+\alpha\overline{\tau}_{2}}\,$, (54)
where $\overline{\tau}_{1}$ and $\overline{\tau}_{2}$ are independent identically distributed random variables with the probability density functions $P(\overline{\tau})$ in eq. (37) for the Model I or in (38) for the Model II. Respectively, the first moment of the splitting probability is given by
[…]
while the second moment obeys
$\begin{split}\Big\langle E^{2}_{-}\Big\rangle&=(1-\alpha)^{2}\int^{\infty}_{0}d\overline{\tau}_{1}\int^{\infty}_{0}d\overline{\tau}_{2}\frac{\overline{\tau}^{2}_{1}P(\overline{\tau}_{1})P(\overline{\tau}_{2})}{\left((1-\alpha)\overline{\tau}_{1}+\alpha\overline{\tau}_{2}\right)^{2}}\,.\end{split}$ (56)
Figure 6: Splitting probabilities for the Models I and II. Left panel: First moment of the splitting probability, eqs. (57), as function of $\alpha$ for $\beta^{2}\sigma^{2}=0$ (grey line, $\langle E_{-}\rangle\equiv 1-\alpha$, no disorder), $\beta^{2}\sigma^{2}=2$ (blue solid curve - Model I and blue dashed curve - Model II) and $\beta^{2}\sigma^{2}=10$ (green solid curve - Model I and green dashed curve - Model II). Right panel: The relative variance $R_{E}=\langle E^{2}_{-}\rangle/\langle E_{-}\rangle^{2}-1$ as function of $\alpha$ for $\beta^{2}\sigma^{2}=2$ (blue curves) and $\beta^{2}\sigma^{2}=10$ (green curves).
Changing the integration variable in the inner integral, $\overline{\tau}_{2}=\overline{\tau}_{1}z$ and performing the integral over $\overline{\tau}_{1}$, we find the following results for the first moment of the splitting probability in both Models,
$\begin{split}\Big\langle E_{-}^{(I)}\Big\rangle&=\frac{1-\alpha}{2\sqrt{\pi\beta^{2}\sigma^{2}}}\int^{\infty}_{0}\frac{dz}{z\left(1-\alpha+\alpha z\right)}\exp\left(-\frac{\ln^{2}\left(z\right)}{4\beta^{2}\sigma^{2}}\right)\,,\\ \Big\langle E_{-}^{(II)}\Big\rangle&=\frac{1-\alpha}{\sqrt{2\pi\beta^{2}\sigma^{2}}}\int^{\infty}_{0}\frac{dz}{z(1-\alpha+\alpha z)}\exp\left(-\frac{\ln^{2}z}{2\beta^{2}\sigma^{2}}\right)\,.\end{split}$ (57)
The integral in eq. (57) cannot be performed exactly but is amenable to asymptotic analysis and, of course, can be evaluated numerically in Mathematica. In the left panel in Fig. 6 we depict the first moment of the splitting probability as function of $\alpha$ together with the trivial result $E_{-}=1-\alpha$ (grey line) which holds for homogeneous systems, i.e., for $\beta\sigma=0$. We observe that upon an increase of disorder, i.e., of the parameter $\beta\sigma$, the first moment of the splitting probability start to deviate from the straight grey line and tends to a constant value $\langle E_{-}\rangle=1/2$ in a progressively larger vicinity of the center of the interval, $\alpha=1/2$. In other words, the value of the disorder-averaged splitting probability becomes less sensitive to the actual starting point. Upon an increase of disorder, behavior of $\langle E_{-}\rangle$ becomes more and more singular in the vicinity of $\alpha=0$ (when $\langle E_{-}\rangle\equiv 1$) and $\alpha=1$ (when $\langle E_{-}\rangle\equiv 0$). Note also that the behavior of $\langle E_{-}\rangle$ is slightly different quantitatively for Models I and II, but qualitatively it is the same.
In a similar way, we calculate the second moment of the splitting probability to get
$\begin{split}\Big\langle\left(E_{-}^{(I)}\right)^{2}\Big\rangle&=\frac{(1-\alpha)^{2}}{2\sqrt{\pi\beta^{2}\sigma^{2}}}\int^{\infty}_{0}\frac{dz}{z\left(1-\alpha+\alpha z\right)^{2}}\exp\left(-\frac{\ln^{2}\left(z\right)}{4\beta^{2}\sigma^{2}}\right)\,,\\ \Big\langle\left(E_{-}^{(II)}\right)^{2}\Big\rangle&=\frac{(1-\alpha)^{2}}{\sqrt{2\pi\beta^{2}\sigma^{2}}}\int^{\infty}_{0}\frac{dz}{z(1-\alpha+\alpha z)^{2}}\exp\left(-\frac{\ln^{2}z}{2\beta^{2}\sigma^{2}}\right)\,.\end{split}$ (58)
Capitalizing on the above expressions we determine the relative variance of the splitting probabilities for the Model I and II, which we depict in the right panel in Fig. 6. We observe that the splitting probability is not self-averaging for both models, except for $\alpha=0$ and apparently for $\alpha=1$.
@@ 5 The realization-dependent mean first-passage times
Consider a finite chain with $N+1$ sites $x$, $x=0,1,\ldots,N$, in which the site $x=N$ is the "target". A particle starts from the site at the left extremity of the chain, i.e., from $x=0$, performs a random walk in a continuous-time between the nearest-neighboring sites of the chain and reaches for the first time the right extremity at some random time instant, which is called the first-passage time to $N$. The mean first-passage time $T_{N}$, with the averaging being performed over all possible realizations of the random walk process for a fixed realization of the transition rates $W_{x,x\pm 1}$, is an important characteristic property of random transport processes which has been amply analyzed in the past (see, e.g., [48, 49, 50] and [65, 66, 67, 68, 69]). Here we follow an early work [70] in which an explicit expression for $T_{N}$ has been derived for the continuous-time and discrete-space master equations for the general form of the transition rates. Adapting this result to our notations, one has

### v3

@@ Abstract
We study random walks evolving in continuous time on a one-dimensional lattice where each site $x$ hosts a quenched random potential $U_{x}$. The potentials on different sites are independent, identically distributed Gaussian random variables. We analyze three distinct models that specify how the transition rates depend on $U_{x}$: the random-force-like model, random walks with randomized stepping times, and the Gaussian trap model. Our analysis focuses on five key disorder-dependent quantities defined for a finite chain with $N$ sites: the probability current, its reciprocal (the resistance), the splitting probability $E_{-}$, the mean first-passage time $T_{N}$, and the diffusion coefficient $D_{N}$ in a periodic chain. By determining the moments of these random variables, we demonstrate that the probability current and resistance are not self-averaging, which leads to pronounced differences between their average and typical behaviors. In contrast, $E_{-}$, $T_{N}$ and $D_{N}$ become self-averaging when $N\to\infty$, though they exhibit strong sample-to-sample fluctuations for finite $N$.
Keywords: Random walks in disordered media, Gaussian random potential, self-averaging vs non-self-averaging
submitted to J. Phys. A: Math. Gen.
@@ 1 Introduction
In this paper, we revisit the long-standing problem of random walks in the presence of a quenched random Gaussian potential (Gaussian white noise), a topic that has received a significant attention in the past (see, e.g., [51]). Specifically, we study dynamics of a particle that undergoes, in continuous time, a random walk between nearest-neighboring sites of a one-dimensional lattice, where each site $x$ hosts a quenched random potential $U_{x}$, which influences the local transition rates. The variables $U_{x}$ are independent and identically distributed (iid) random variables, with a common Gaussian distribution. This model can be also viewed from a somewhat different perspective. Namely, it can be considered as an effective description, in terms of the Fick-Jacob’s approximation (see, e.g., [60, 61]), of diffusion in a narrow two- or three-dimensional channel with Gaussian-distributed random wall corrugation.
@@
While the statistics of the last two quantities have been explored in previous studies and some well-established results have been obtained (see, e. g., [49, 48, 51]), a comprehensive analysis of the first three variables is lacking at present. It is also important to note that all the random variables under study are formally defined in the limit $t\to\infty$. Consequently, our analysis does not address transient effects, such as aging and the ensuing sub-diffusive behavior, which have been extensively discussed in the literature on the random force and random traps models (see, e. g., [49]).
@@
The paper is outlined as follows: In Sec. 2 we present our model and describe three different dynamical scenarios. In Sec. 3 we concentrate on the statistical properties of the realization-dependent currents through a finite chain and of its reciptocal value - the resistance of a finite chain with respect to a random transport. In Sec. 4 we discuss the behavior of the realization-dependent splitting probabilities. Section 5 is devoted to the analysis of the mean first-passage times through finite intervals. Next, in Sec. 6 we evaluate the moments of the diffusion coefficient in period chains in the limit $N\to\infty$ for all three dynamical scenarios. In Sec. 7 we conclude with a brief recapitulation of our results. Some derivations are relegated to Appendices. In A we present the derivation of the probability currents through a finite chain with fixed disorder for all three dynamical scenarios, while in B we express the realization-dependent splitting probability through the resistances of the intervals from the right and from the left of the starting point $x=x_{0}$. Lastly, in C, capitalizing on the exact result due to Derrida [4], we derive exact compact expressions for the realization-dependent diffusion coefficient on a periodic chain for all three dynamical scenarios.
@@ 2.2 Scenario II: Random walk with randomized stepping-times
Recalling that each site of the lattice hosts a random potential $U_{x}$, we define the transition probabilities $p_{x,x\pm 1}$ in the way which resembles the definition of the transition rates in the above Model I, except that these probabilities are normalized to unity. That being, we stipulate that
@@ 3.1 The moments of resistances
The above Eqs. (28) and (29) imply that the resistances in Models I and II are not self-averaging in the limit $N\to\infty$. The resistance in the Model III is not self-averaging for any value of $N$. Moreover, in all three Models the relative variances $R_{\tau}$ show an unbounded growth with $\beta\sigma$, which signifies that fluctuations become more important the larger the disorder is. The behavior of $R_{\tau}$ for all three Models is depicted in Fig. 2. Below we will show that such a lack of self-averaging is essentially the boundary effect.
Figure 2: The relative variances $R_{\tau}$ of the resistances $\tau_{N}$ in the Models I (left panel), II (middle panel) and III (right panel) as functions of $\beta\sigma$. Solid curves are our analytical predictions in Eqs. (28) to (30) for $N=\infty$, while the symbols depict $R_{\tau}$ evaluated by a numerical averaging of $\tau_{N}$ and $\tau^{2}_{N}$ for finite values of $N$ (see the insets).
@@
It is easy to find that the first two moments of the latter probability density function are given precisely by Eq. (LABEL:momtau2). Note, as well, that the probability density functions in Eqs. (37) and (38) have the form of the log-normal distribution as the parental form in Eq. (5), and differ from it (and between themselves) only in the pre-exponential algebraic dependence on the reduced resistance and by the normalizations.
@@
– The lack of self-averaging is solely due to fluctuations of the on-site potentials at the site $x=0$ (for Models I and III) and at the sites $x=0,1$ (for Model II). The contributions of the sites in the bulk of the chain are self-averaging for Models I and II, and is constant, i.e., non-fluctuating for the Model III.
@@ 3.3 Probability currents
i.e., the typical current in these two Models is characterized by the same decreasing, super-Arrhenius function of disorder and the temperature and therefore, is much smaller than the averaged currents in eqs. (43). Such a strong "reduction" of the current is apparently due to the fact that in typical realizations of disorder a particle will always get localized between two sites at which sufficiently large forces point in the opposite directions. Further on, we find that in the Model III the typical current obeys
@@ 3.4 Sample-to-sample fluctuations of the resistances and currents
This probability density function is depicted in Fig. 5 for three values of the parameter $\beta\sigma$. We observe that for a relatively small disorder, i.e., for $\beta\sigma=0.1$, $P(\omega)$ is unimodal and centered at $\omega=1/2$. This signifies that for a small disorder it is indeed most likely that the values of resistances (or currents) observed in two randomly chosen samples are very close to each other. Upon an increase of $\beta\sigma$, for $\beta\sigma=1$, we encounter a different behavior when the central part of $P(\omega)$ becomes flat and $\omega=1/2$ is no longer the most probable value. In this situation, (almost) any relation between the two resistance becomes equally probable. Increasing $\beta\sigma$ further, we have that already for $\beta\sigma=1.2$ the probability density function becomes bimodal with two maxima close to $\omega=0$ and $\omega=1$. Here, $\omega=1/2$ corresponds not to the most probable but to the least probable value. This signifies that the resistance (or currents) observed in two sample are most likely disproportionally different and sample-to-sample fluctuations are very significant.
@@ 4 The realization-dependent splitting probability
Consider a one-dimensional chain composed of $N+1$ sites labeled $x=0,1,\ldots,N$. Suppose that the particle starts at site $x=x_{0}$ and performs a continuous-time random walk with site-dependent transition rates $W_{x,\,x+1}$ (to the right, $x\to x+1$) and $W_{x,\,x-1}$ (to the left, $x\to x-1$). The sites $x=0$ and $x=N$ at the boundaries are absorbing, meaning that once the particle reaches either boundary, it remains there forever. We focus here on the so-called splitting probability $E_{-}(x_{0})$; that is, the probability that a particle starting from the site $x_{0}$ is eventually absorbed at the left boundary $x=0$ before ever reaching the right boundary $x=N$.
In B we show that the realization-dependent splitting probability admits the following compact representation,
$E_{-}(x_{0})=\frac{\displaystyle\tau_{N-x_{0}}^{+}}{\displaystyle\tau_{x_{0}}^{-}+\tau_{N-x_{0}}^{+}},$ (53)
where $\tau_{x_{0}}^{+}$ and $\tau_{N-{x_{0}}}^{-}$ are the resistances of the intervals $(0,x_{0}-1)$ and $(x_{0}+1,N)$ from the left and from the right of the starting point $x_{0}$, respectively. These realization-dependent resistances are defined explicitly by rather lengthy expressions (see Eqs. (107)) which we present in B. An analogous expression for the continuous-space system can be found in [45].
As shown in B, the realization-dependent splitting probability for Model III is trivially $E_{-}(x_{0})=1-x_{0}/N$, i.e., it is not a fluctuating property and is the same as in homogeneous systems without disorder [66]. Concurrently, for the Models I and II the splitting probabilities do depend on disorder and for both Models are given by
$E_{-}(x_{0})=\frac{H_{x_{0},\,N}}{H_{0,N}}\,$, (54)
where the function $H_{x_{0},\,N}$ is defined by Eq. (23) with the first index replaced by $x_{0}$. A straightforward analysis shows that in the joint limit $N,\,x_{0}\to\infty$ with the ratio $\alpha=x_{0}/N$ kept fixed,
$\frac{1}{N}\sum_{x=0}^{N-1}\phi_{x}\phi_{x+1}\;\xrightarrow[N\to\infty]{}\;\langle\phi\rangle^{2},\qquad\frac{1}{N-x_{0}}\sum_{x=x_{0}}^{N-1}\phi_{x}\phi_{x+1}\;\xrightarrow[N,x_{0}\to\infty]{}\;\langle\phi\rangle^{2},$
almost surely. As a consequence, the realization-dependent splitting probability self-averages in this limit and converges to
$E_{-}(x_{0})\simeq 1-\frac{x_{0}}{N},$ (55)
with vanishing sample-to-sample fluctuations.
Figure 6: The splitting probability $E_{-}(x_{0})$ as function of $\alpha=x_{0}/N$. Top row: The splitting probability in Eq. (54) for a given realization of disorder on a chain with $N=10^{5}$ sites for different temperatures (left panel). Numerically-evaluated first moment of the splitting probability for several values of $N$ at fixed inverse temperature $\beta=10$ (right panel). In both panels the dashed line depicts the analytical prediction in Eq. (55). Bottom row: Relative variance $R_{E}=\mathrm{Var}(E_{-}(x_{0}))/\langle E_{-}(x_{0})\rangle^{2}$ for $\beta=2$ and several values of $N$.
In Fig. 6 we show the splitting probability as a function of the scaled position $\alpha=x_{0}/N$. The top row (left panel) displays the realization-dependent $E_{-}(x_{0})$ given by Eq. (53), evaluated for a single randomly generated disorder realization on a chain of $10^{5}$ sites, at several temperatures. At the lowest temperature ($T=0.01$), the sum $H_{0,N}$ is entirely dominated by its largest term. As a result, the splitting probability takes the form of a step function: $E_{-}(x_{0})=1$ up to a critical realization-dependent value of $\alpha$, beyond which it drops abruptly to zero. When the temperature is increased by an order of magnitude, subdominant contributions to $H_{0,N}$ become relevant. Consequently, $E_{-}(x_{0})$ develops a descending staircase-like structure as a function of $\alpha$, with each plateau corresponding to dominance of of different terms in the sum. Upon further increase of the temperature, the realization-dependent fluctuations are progressively smoothed out. In this regime, the realization-dependent splitting probability becomes nearly indistinguishable from that of a homogeneous system, recovering the linear behavior $E_{-}(x_{0})=1-\alpha$ [66]. Next, the top-right panel of Fig. 6 shows the numerically evaluated disorder-averaged splitting probability for several values of $N$ at a fixed temperature. We find that all curves are nearly indistinguishable from the asymptotic prediction in Eq. (55), indicating that size-dependent corrections remain very small, even at the relatively low temperature $T=0.1$. Lastly, the bottom row of Fig. 6 displays the relative variance of the realization-dependent splitting probability as a function of $\alpha=x_{0}/N$ for several values of $N$ (see inset), at a fixed inverse temperature $\beta=2$. Away from the boundaries $\alpha=0$ and $\alpha=1$, where $E_{-}$ equals $1$ and 0, respectively, and hence $R_{E}\equiv 0$, the relative variance is nonzero for finite $N$ but decreases systematically as $N$ increases. In the limit $N\to\infty$, $R_{E}$ vanishes, indicating that the splitting probability becomes self-averaging.
@@ 5 The realization-dependent mean first-passage times
Consider a finite chain with $N+1$ sites $x$, $x=0,1,\ldots,N$, in which the site $x=N$ is the "target". A particle starts from the site at the left extremity of the chain, i.e., from $x=0$, performs a random walk in a continuous-time between the nearest-neighboring sites of the chain and reaches for the first time the right extremity at some random time instant, which is called the first-passage time to $N$. The mean first-passage time $T_{N}$, with the averaging being performed over all possible realizations of the random walk process for a fixed realization of the transition rates $W_{x,x\pm 1}$, is an important characteristic property of random transport processes which has been amply analyzed in the past (see, e.g., [48, 49, 50] and [65, 66, 67, 68, 69]). Here we follow an early work [70] in which an explicit expression for $T_{N}$ has been derived for the continuous-time and discrete-space master equation for the general form of the transition rates. Adapting this result to our notations, one has